# AdvAgent: Controllable Blackbox Red-teaming on Web Agents

### v3

@@ Abstract
Foundation model-based agents are increasingly used to automate complex tasks, enhancing efficiency and productivity. However, their access to sensitive resources and autonomous decision-making also introduce significant security risks, where successful attacks could lead to severe consequences. To systematically uncover these vulnerabilities, we propose AdvAgent, a black-box red-teaming framework for attacking web agents. Unlike existing approaches, AdvAgent employs a reinforcement learning-based pipeline to train an adversarial prompter model that optimizes adversarial prompts using feedback from the black-box agent. With careful attack design, these prompts effectively exploit agent weaknesses while maintaining stealthiness and controllability. Extensive evaluations demonstrate that AdvAgent achieves high success rates against state-of-the-art GPT-4-based web agents across diverse web tasks. Furthermore, we find that existing prompt-based defenses provide only limited protection, leaving agents vulnerable to our framework. These findings highlight critical vulnerabilities in current web agents and emphasize the urgent need for stronger defense mechanisms. We release code at https://ai-secure.github.io/AdvAgent/.
@@ 1 Introduction
Recent efforts have introduced adversarial attacks against generalist web agents to expose vulnerabilities prior to real-world deployment (Yang et al., 2024; Wang et al., 2024). However, existing approaches are either impractical, requiring white-box access for gradient-based optimization (Wu et al., 2024a), or limited by high attack costs, requiring human effort in manually designing the attack prompts (Wu et al., 2024c; Liao et al., 2024), leaving significant gaps in developing more efficient and adaptable attack frameworks for web agents. While adversarial attacks have been proposed for LLMs and VLMs to optimize attack prompts automatically (Guo et al., 2024; Huang et al., 2024), these methods lack the flexibility to target LLM and VLM agents (Zou et al., 2023) and struggle to achieve effective transferability in black-box attack settings (Liu et al., 2024c).
@@ 4 AdvAgent: Controllable Black-box Attacks on Web Agents
Figure 2: AdvAgent Prompter Model Training. During data collection, we first collect the training dataset using LLM-based attack prompter by Algorithm 2 in Appendix A. Then we collect positive and negative feedback from the target black-box model. During prompter model training, we first launch the first stage SFT using the positive subsets. The model is further trained in the second DPO stage using both positive and negative feedback.
@@ 4.1 Automatic Attack and Feedback Collection
Automatic Attack and Feedback Collection Pipeline. Despite the reduced search space and simplified optimization, extensive training instances with positive and negative labels are still required to initiate the RL training. To ensure the diversity of the training instances, we employ LLMs as an attack prompter, generating a set of $n$ various diverse adversarial prompts $\{q_{i}\}_{i=1}^{n}$, as illustrated in Algorithm 2 in Appendix A. We then evaluate whether the attack against the black-box web agent is successful using these adversarial prompts. Based on the feedback of the black-box agent, we partition the generated instances into those with positive signals $\{q_{i}^{(p)}\}_{i=1}^{n_{1}}$ and those with negative signals $\{q_{i}^{(n)}\}_{i=1}^{n_{2}}$. These partitions are subsequently used for RL training. The process is illustrated in Figure 2. We also show pairs of adversarial prompts with subtle differences but different attack results in Figure 4.
@@ 4.2 Adversarial Prompter Model Training
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{h}\sum_{i=1}^{n_{1}}\log\pi_{\theta}(q_{i}^{(p)}|h)$ (2)
@@ 5.3 In depth analysis of AdvAgent
AdvAgent demonstrates high controllability for targeting different attack goals. We evaluate the controllability of AdvAgent by modifying the attack targets of successful adversarial injections to previously unseen targets. As shown in Table 3, AdvAgent achieves an average ASR of $98.5\%$ across different domains for new targets, with additional results for the Gemini 1.5-powered agent provided in Table 6 (see Appendix B). These results confirm that AdvAgent’s adversarial injections are not only highly effective but also easily controllable, allowing attackers to switch targets with minimal effort and no additional computational overhead.
@@ 7 Conclusion
To uncover the vulnerabilities of web agents in real-world scenarios, we propose AdvAgent, a black-box targeted red-teaming framework designed to evaluate web agents across various domains and tasks. Extensive experiments demonstrate that AdvAgent consistently achieves significantly higher ASRs than existing baselines, effectively compromising web agents powered by different proprietary backend models. Our findings also reveal that existing prompt-based defenses provide only limited protection against AdvAgent. Despite considering common mitigation strategies, web agents remain highly vulnerable, with ASRs exceeding $88.8\%$ even after applying defenses. This highlights the urgent need for more robust security measures to protect against adversarial attacks. Despite some limitations as we discuss in Appendix E, such as requiring offline feedback for prompt optimization and focusing on step-based ASRs due to current constraints of web agents, our study highlights the critical need for stronger defenses in this domain. By exposing these vulnerabilities through sophisticated red-teaming techniques, we aim to inspire further research into developing effective countermeasures that enhance the security and resilience of web agents.
@@ Acknowledgements
Appendix A Additional Details of AdvAgent
We show the LLM-based Attack Prompter in Algorithm 2, where we leverage LLM to help us generate diverse attack prompts.
Algorithm 2 LLM-based Attack Prompter
0:  Original HTML contents $h$, large-language model LLM
1:  function Prompter$(h,\texttt{LLM})$
[…]
5:  end function
Appendix B Additional Experiment Result
@@
Appendix C Additional Examples
@@
Appendix D Additional Related Work

### v4

@@ Abstract
Foundation model-based agents are increasingly used to automate complex tasks, enhancing efficiency and productivity. However, their access to sensitive resources and autonomous decision-making also introduce significant security risks, where successful attacks could lead to severe consequences. To systematically uncover these vulnerabilities, we propose AdvAgent, a black-box red-teaming framework for attacking web agents. Unlike existing approaches, AdvAgent employs a reinforcement learning-based pipeline to train an adversarial prompter model that optimizes adversarial prompts using feedback from the black-box agent. With careful attack design, these prompts effectively exploit agent weaknesses while maintaining stealthiness and controllability. Extensive evaluations demonstrate that AdvAgent achieves high success rates against state-of-the-art GPT-4-based web agents across diverse web tasks. Furthermore, we find that existing prompt-based defenses provide only limited protection, leaving agents vulnerable to our framework. These findings highlight critical vulnerabilities in current web agents and emphasize the urgent need for stronger defense mechanisms. We release our code at https://ai-secure.github.io/AdvAgent/.
@@ 1 Introduction
Recent efforts have introduced adversarial attacks against generalist web agents to expose vulnerabilities prior to real-world deployment (Yang et al., 2024; Wang et al., 2024; Zhang et al., 2025). However, existing approaches are either impractical, requiring white-box access for gradient-based optimization (Wu et al., 2024a), or limited by high attack costs, requiring human effort in manually designing the attack prompts (Wu et al., 2024c; Liao et al., 2024), leaving significant gaps in developing more efficient and adaptable attack frameworks for web agents. While adversarial attacks have been proposed for LLMs and VLMs to optimize attack prompts automatically (Guo et al., 2024; Huang et al., 2024), these methods lack the flexibility to target LLM and VLM agents (Zou et al., 2023) and struggle to achieve effective transferability in black-box settings (Liu et al., 2024c).
@@ 4 AdvAgent: Controllable Black-box Attacks on Web Agents
Figure 2: AdvAgent Prompter Model Training. During data collection, we first collect the training dataset using LLM-based attack prompter by Algorithm 1 in LABEL:sec:additional_algo. Then we collect positive and negative feedback from the target black-box model. During prompter model training, we first launch the first stage SFT using the positive subsets. The model is further trained in the second DPO stage using both positive and negative feedback.
@@ 4.1 Automatic Attack and Feedback Collection
Algorithm 1 LLM-based Attack Prompter
0:  Original HTML contents $h$, large-language model LLM
1:  function Prompter$(h,\texttt{LLM})$
[…]
5:  end function
Automatic Attack and Feedback Collection Pipeline. Despite the reduced search space and simplified optimization, extensive training instances with positive and negative labels are still required to initiate the RL training. To ensure the diversity of the training instances, we employ LLMs as an attack prompter, generating a set of $n$ various diverse adversarial prompts $\{q_{i}\}_{i=1}^{n}$, as illustrated in Algorithm 1. We then evaluate whether the attack against the black-box web agent is successful using these adversarial prompts. Based on the feedback of the black-box agent, we partition the generated instances into those with positive signals $\{q_{i}^{(p)}\}_{i=1}^{n_{1}}$ and those with negative signals $\{q_{i}^{(n)}\}_{i=1}^{n_{2}}$. These partitions are subsequently used for RL training. The process is illustrated in Figure 2. We also show pairs of adversarial prompts with subtle differences but different attack results in Figure 4.
@@ 4.2 Adversarial Prompter Model Training
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{h}\sum_{i=1}^{n_{1}}\log\pi_{\theta}(q_{i}^{(p)}|h)\vspace{-2pt}$ (2)
@@ 5.3 In depth analysis of AdvAgent
AdvAgent demonstrates high controllability for targeting different attack goals. We evaluate the controllability of AdvAgent by modifying the attack targets of successful adversarial injections to previously unseen targets. As shown in Table 3, AdvAgent achieves an average ASR of $98.5\%$ across different domains for new targets, with additional results for the Gemini 1.5-powered agent provided in Table 6 (see Appendix A). These results confirm that AdvAgent’s adversarial injections are not only highly effective but also easily controllable, allowing attackers to switch targets with minimal effort and no additional computational overhead.
@@ 7 Conclusion
To uncover the vulnerabilities of web agents in real-world scenarios, we propose AdvAgent, a black-box targeted red-teaming framework designed to evaluate web agents across various domains and tasks. Extensive experiments demonstrate that AdvAgent consistently achieves significantly higher ASRs than existing baselines, effectively compromising web agents powered by different proprietary backend models. Our findings also reveal that existing prompt-based defenses provide only limited protection against AdvAgent. Despite considering common mitigation strategies, web agents remain highly vulnerable, with ASRs exceeding $88.8\%$ even after applying defenses. This highlights the urgent need for more robust security measures to protect against adversarial attacks. Despite some limitations as we discuss in Appendix D, such as requiring offline feedback for prompt optimization and focusing on step-based ASRs due to current constraints of web agents, our study highlights the critical need for stronger defenses in this domain. By exposing these vulnerabilities through sophisticated red-teaming techniques, we aim to inspire further research into developing effective countermeasures that enhance the security and resilience of web agents.
@@ Acknowledgements
Appendix A Additional Experiment Result
@@
Appendix B Additional Examples
@@
Appendix C Additional Related Work